OpenStudy (anonymous):

There is a number that is 8 less than the sum of 20 + 12. Find the number. Let n = the number.

OpenStudy (anonymous):

8 less than 20+12 -(8) + 20+12 =24

OpenStudy (anonymous):

32-8 =24

OpenStudy (anonymous):

24

OpenStudy (anonymous):

answer choices are n=8(20+12),n=(20+12)-8,n-8=(20+12),n=(20+12)+8

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

A father is 1 year older than 3 times the age of his son. The son is 20 years old. Let F = the age of the father. Which formula would calculate the father's age? F + 1 = 3(20) F - 1 = 3(20) F = 3(20 - 1)

OpenStudy (anonymous):

3(20)+1

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

second

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Determine which operation should be performed on each side of the equation below to solve for the variable. x + 9 = 20 Add (-9) to both sides. Add (-20) to both sides. Add (+9) to both sides. Add (-29) to both sides.

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

Determine which operation should be performed on each side of the equation below to solve for the variable. -13 = 4 + m Add (+13) to both sides. Add (+4) to both sides. Add (-13) to both sides. Add (-4) to both sides.

OpenStudy (nirmalnema):

add-4 both side

OpenStudy (anonymous):

fourth

OpenStudy (anonymous):

Determine which operation should be performed on each side of the equation below to solve for the variable. x - 8 = 2 Add 8 to both sides. Add (-2) to both sides. Add 2 to both sides. Add (-8) to both sides

OpenStudy (nirmalnema):

1st one is correct

OpenStudy (anonymous):

Determine which operation should be performed on each side of the equation below to solve for the variable. 9 = -3 + r Add 3 to both sides. Add 9 to both sides. Add 6 to both sides. Add (-3) to both sides.

OpenStudy (nirmalnema):

add 3 to both side

OpenStudy (anonymous):

first

OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? p + (-2) = 21

OpenStudy (nirmalnema):

2 is to be added

OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? 18 + y = 32

OpenStudy (nirmalnema):

-18

OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? 18 + y = 32

OpenStudy (nirmalnema):

same as above.